OpenStudy (anonymous):

solve for θ in the following equation:Sin 2θ=cosθ

OpenStudy (anonymous):

^^

OpenStudy (jamesj):

I'll help you quickly this:

OpenStudy (anonymous):

no3 question haha thank u

OpenStudy (jamesj):

You should have a formula for sin(2.theta) \[\sin(2\theta) = 2 \sin \theta . \cos \theta\]

OpenStudy (jamesj):

Hence if sin(2.theta) = cos(theta), that implies .... what? Simplify the expression.

OpenStudy (anonymous):

u said before 1 right?

OpenStudy (jamesj):

?

OpenStudy (anonymous):

w8

OpenStudy (anonymous):

if i simply it with my poor brain i think they are the same?

OpenStudy (jamesj):

Sin 2θ=cosθ => 2 sin θ . cosθ = cosθ => 2 sin θ = 1 => sin θ = 1/2

OpenStudy (jamesj):

Now you just have to say for what values of θ is sin θ = 1/2

OpenStudy (anonymous):

1/2

OpenStudy (jamesj):

What is sin(30) and sin(60)?

OpenStudy (anonymous):

sin(30)

OpenStudy (jamesj):

(You should learn by heart the sin, cos and tan all these angles 0, 30, 45, 60, 90.)

OpenStudy (jamesj):

What is the value of sin(30 degrees)?

OpenStudy (anonymous):

1/2

OpenStudy (jamesj):

right. Now are there other θ for which sin θ = 1/2. Yes. So you need to write them all down.

OpenStudy (anonymous):

ok